# Compressed and Smooth Latent Space for Text Diffusion Modeling

### v1

@@ Results.
[…]
@@ Appendix A Dataset descriptions
Appendix B Implementation details
B.1 Latent diffusion pipeline
With a diffusable latent manifold in hand (§4.2.2), we now describe the diffusion model that operates entirely within this space. Our setup follows Eq. (2): we train a denoiser that receives a noisy latent $\mathbf{z}_{t}$ at diffusion time $t\!\in\![0,1]$ and predicts the corresponding clean latent $\mathbf{z}_{0}$. The loss is the simple mean-squared error $\lVert\mathbf{z}_{0}-z_{\theta}(z_{t},t)\rVert_{2}^{2}$, where $z_{\theta}$ shares parameters across all time-steps. In line with earlier work [18, 33], we employ self-conditioning: on $50\%$ of training updates the model is fed its own previous estimate of $\mathbf{z}_{0}$ as an additional input. This iterative refinement enables the model, at inference time, to reuse its past predictions, producing markedly crisper and more coherent text. We adopt the noise schedule introduced by TEncDM [33]: $\alpha_{t}=\frac{1}{1+\tan(t\pi/2)^{2}\cdot d^{2}}$, where parameter $d$ controls he rate at which noise is injected during the diffusion steps. At inference time we use Euler solver with 200 uniform steps. The denoiser itself inherits the TEncDM [33] architecture. It is a 12-layer Transformer with 12 attention heads and a hidden size of 768. In total, the network comprises roughly $\approx\!130$M parameters.
@@ B.1 Latent diffusion pipeline
B.2 Training and inference configuration
[…]
@@ B.2 Training and inference configuration
B.3 Inference time benchmarking configuration
@@
Appendix C Additional autoencoder training analysis
C.1 Marginal utility of variational prior in text autoencoder
@@
where $\mathrm{CE}(w,\hat{w})$ is the token-level cross-entropy, $\mathrm{MSE}(h,\hat{h})$ matches reconstructed contextual representations, and $\beta\!\geq\!0$ balances reconstruction fidelity against latent regularisation.
@@
In our experiments, we observed that the performance is highly sensitive to the choice of  $\beta$. Table 6 presents the results for different values of  $\beta$.
@@
This limited improvement can be attributed to the decoder’s robustness to Gaussian noise in the latent space. As illustrated in Figure 4, the decoder maintains reasonable performance even when sampling from the prior with $\sigma=1$, indicating that the latent space is sufficiently robust without strong KL regularization.
@@
C.2 Empirical analysis of autoencoder training regimes on Wikipedia dataset
Table 7: Comparison of autoencoder training regimes on text generation quality on Wikipedia. Features are added cumulatively from top to bottom.
Configuration MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
CE($\bf{w,\hat{w}}$) $0.056$ $385.6$ $0.590$
+ MSE($\bf{h,\hat{h}}$) $0.069$ $344.4$ $0.625$
+ Random masking (rate)
   0.3 $0.081$ $289.1$ $0.594$
+ Gaussian noising ($\delta$)
   0.7 $0.103$ $205.0$ $0.588$
+ Latent dropout (rate)
   0.4 $0.112$ $186.7$ $0.581$
We also conduct a large-scale ablation on the Wikipedia dataset with $N=16$ latents and max sequence length $L=128$, following a setup analogous to that in section 6.1. Results, presented in Table 7, reveal trends consistent with those observed in the small-scale ablation. Notably, applying the proposed training modifications leads to substantial improvements in evaluation metrics. The final model approaches the performance of the TEncDM baseline [33], while operating in a latent space that is 8×8$\times$8 × smaller.
Additionally, we investigate how scaling the latent space length on Wikipedia affects performance. We experiment with latent sequence lengths ranging from $16$ to $128$. As expected, we observe steady performance improvements with longer latent sequences. Results are provided in Table 8.
Table 8: Impact of scaling the number of latent vectors $N$ on unconditional generation quality on Wikipedia.
$N$ MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
Source 0.953 21.7 0.403
BERT repr. 0.109 173.1 0.562
16 0.112 186.7 0.581
32 0.149 134.4 0.579
64 0.157 128.1 0.582
128 0.173 118.2 0.592
C.3 Additional details of comparison across generative paradigms
@@
We also present extended results for our conditional generation tasks using a broader variety of evaluation metrics. For XSum and SQuAD2.0, we report ROUGE scores [15], a standard metric that evaluates text quality based on n-gram overlap with reference outputs. For ParaDetox, we additionally include the J-score, which is defined as the product of style accuracy, fluency, and content preservation. Extended results are summarized in Table 9.
Appendix D Limitations
@@
Appendix E Societal Impact
@@
Appendix F Generation examples

### v2

@@ Results.
As shown in Table˜9, CosmosN=16N=16 achieves strong performance relative to its closest latent diffusion baseline, TEncDM. It consistently outperforms TEncDM across most evaluation metrics, with the exception of unconditional perplexity, while offering significantly faster generation. Scaling up to a bigger latent representation, CosmosN=128N=128 further advances generation quality. On ROCStories, it achieves a MAUVE score of $0.940$, closely approaching the human reference score of $0.953$, and substantially outperforming the best autoregressive baseline (GPT2), which reaches only $0.789$. Although CosmosN=128N=128 lags behind GPT2 in perplexity, this gap reflects an evaluation bias: perplexity inherently favors autoregressive models, particularly when assessed using the same decoding objective they were trained on.
In generation tasks that involve longer input contexts, CosmosN=128N=128 consistently matches or slightly exceeds the performance of both diffusion-based and autoregressive baselines, while substantially reducing inference time. This efficiency stems from a core advantage of latent diffusion: unlike autoregressive decoding, where inference time grows linearly with sequence length, diffusion models operate with a fixed number of sampling steps. Although this cost is less favorable for short sequences, the benefits become increasingly pronounced as input length grows. This favorable scaling behavior makes latent diffusion as a promising direction for future research.
8 Experiments on OpenWebText
In this section, we present an empirical evaluation of our proposed method on the large-scale OpenWebText (OWT) (Gokaslan2019OpenWeb, ) dataset.
8.1 Impact of Diffusion Model Depth
Table 5: Scaling results on OpenWebText for 128-token generation. Increasing the number of layers in the diffusion model transformer improves generation quality and diversity. Best results in bold.
Model Size MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$ Mem $\downarrow$
0.12 B 0.849 97.6 0.492 0.135
0.25 B 0.914 91.2 0.546 0.124
0.5 B 0.923 89.7 0.554 0.125
We train three diffusion models with varying depths of 12, 24, and 48 layers, which correspond to approximately 0.12B, 0.25B, and 0.5B parameters, respectively. For these experiments, the autoencoder and latent space configuration are held constant. We selected our best-performing autoencoder from preliminary experiments, which utilizes the largest latent space of 128 vectors for 128-token sequences.
The results presented in Table 5 clearly indicate a positive scaling trend. As the number of layers in the diffusion model increases, both generation quality and diversity improve. This demonstrates that increasing the capacity of the diffusion component is an effective strategy for enhancing performance, even when the autoencoder’s architecture remains fixed.
8.2 Comparison on Long-Sequence Generation
Table 6: Comparison with the GIDD masked diffusion model on OpenWebText for 512-token generation. Our model (CosmosN=512) shows significantly better quality and coherence.
Model MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$ Mem $\downarrow$
Source text 0.968 23.2 0.465 0.036
GIDD 0.286 228.3 0.588 0.112
TEncDM 0.228 118.6 0.324 0.102
CosmosN=256 0.263 84.3 0.368 0.156
CosmosN=512 0.702 55.0 0.319 0.180
We further assess our model’s capabilities on the challenging task of generating 512-token sequences. We compare our model, Cosmos, in two configurations: without compression ($N=512$ latent vectors) and with 2x compression ($N=256$). These are benchmarked against TEncDM (shabalin2025tencdm, ), a latent diffusion baseline, and GIDD (von2025generalized, ), a strong discrete diffusion model. The results, presented in Table 6, offer several key insights into the advantages of our approach.
First, our primary model, CosmosN=512, demonstrates a significant performance gap over the discrete diffusion paradigm. It outperforms GIDD by a large margin, achieving a MAUVE score that is nearly three times higher (0.792 vs. 0.286) and a perplexity that is almost four times lower (55.0 vs. 228.3), indicating superior text quality and coherence.
Second, the results confirm the expected trade-off between compression and generation fidelity. While our compressed model, CosmosN=256, shows a degradation in performance compared to its uncompressed counterpart, it still outperforms TEncDM, a baseline that works with full-length, uncompressed textual representations. This result powerfully highlights the efficiency of our latent diffusion framework for text generation.
@@ 9 Conclusion
Acknowledgements
The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4E0002 and the agreement with HSE University № 139-15-2025-009. This research was supported in part through computational resources of HPC facilities at HSE University.
@@ Appendix A Dataset descriptions
Appendix B Additional Related Work
In addition to the works discussed in the main paper, a significant body of research has explored diffusion models for text generation. These can be broadly categorized as follows.
B.1 Gaussian Diffusion on Embeddings
One line of work focuses on applying Gaussian diffusion directly to word embeddings. Models like DiffuSeq [10] and SeqDiffuSeq [45] are designed for sequence-to-sequence text generation tasks. DiffuSeq employs a partial noising strategy and is trained end-to-end in a classifier-free manner. SeqDiffuSeq utilizes an encoder-decoder Transformer architecture and incorporates self-conditioning and an adaptive noise schedule to enhance performance. AR-Diffusion [43] integrates autoregressive principles into the diffusion process to better capture the sequential dependencies in language, allowing tokens on the right to depend on the generated tokens on the left. This is achieved by using a dynamic number of denoising steps that varies with the token’s position.
B.2 Simplex-Based Diffusion
Another direction explores alternative spaces for the diffusion process. TESS [20], for instance, applies a fully non-autoregressive diffusion process on the logit simplex space. This approach, combined with a novel self-conditioning technique, demonstrates strong performance on various natural language generation tasks while requiring fewer diffusion steps.
B.3 Masked Diffusion
Masked-token diffusion models adopt a BERT-style prediction objective with iterative unmasking, enabling fast and parallel generation. For natural language, SEDD [18] established a connection between score-based models and ELBO maximization, demonstrating the effectiveness of the approach. Subsequent work, such as simple absorbing-mask diffusion with a clean training recipe, has successfully narrowed the perplexity gap to autoregressive language models and supports flexible semi-autoregressive decoding [34]. The quality of generation is also significantly boosted by various inference-time techniques. These include remasking for iterative correction [41], discrete guidance for improved controllability [35], and hybrid methods that expand self-correction capabilities [40].
B.4 Latent Diffusion
In the realm of latent diffusion, several alternative baseline models have been proposed. LD4LG [19] learns a continuous diffusion model in the latent space of a pretrained encoder-decoder model. This method samples continuous latent representations which are then decoded into natural language, proving effective for both unconditional and conditional text generation. Similarly, PLANNER [47] adopts a two-stage recipe where a fine-tuned BERT encoder produces a variational latent code, which is then refined by a diffusion model before being translated into text by a GPT-2 decoder. TEncDM [36], on the other hand, operates in the space of pretrained language model encodings, which, unlike embeddings, integrate contextual information.
Appendix C Implementation details
C.1 Latent diffusion pipeline
With a diffusable latent manifold in hand (§4.2.2), we now describe the diffusion model that operates entirely within this space. Our setup follows Eq. (2): we train a denoiser that receives a noisy latent $\mathbf{z}_{t}$ at diffusion time $t\!\in\![0,1]$ and predicts the corresponding clean latent $\mathbf{z}_{0}$. The loss is the simple mean‑squared error $\lVert\mathbf{z}_{0}-z_{\theta}(z_{t},t)\rVert_{2}^{2}$, where $z_{\theta}$ shares parameters across all time‑steps. In line with earlier work [19, 36], we employ self‑conditioning: on $50\%$ of training updates the model is fed its own previous estimate of $\mathbf{z}_{0}$ as an additional input. This iterative refinement enables the model, at inference time, to reuse its past predictions, producing markedly crisper and more coherent text. We adopt the noise schedule introduced by TEncDM [36]: $\alpha_{t}=\frac{1}{1+\tan(t\pi/2)^{2}\cdot d^{2}}$, where parameter $d$ controls he rate at which noise is injected during the diffusion steps. At inference time we use Euler solver with 200 uniform steps. The denoiser itself inherits the TEncDM [36] architecture. It is a 12‑layer Transformer with 12 attention heads and a hidden size of 768. In total, the network comprises roughly $\approx\!130$M parameters.
@@ C.1 Latent diffusion pipeline
C.2 Training and inference configuration
All models are trained on 8 NVIDIA A100 GPUs. Detailed training configurations and approximate durations for both CosmosN=128N=128 and CosmosN=16N=16 are provided in Table 7.
@@ C.2 Training and inference configuration
C.3 Inference time benchmarking configuration
@@
Appendix D Additional autoencoder training analysis
D.1 Marginal utility of variational prior in text autoencoder
@@
[…]
@@
In our experiments, we observed that the performance is highly sensitive to the choice of  $\beta$. Table˜8 presents the results for different values of  $\beta$.
@@
This limited improvement can be attributed to the decoder’s robustness to Gaussian noise in the latent space. As illustrated in Figure˜4, the decoder maintains reasonable performance even when sampling from the prior with $\sigma=1$, indicating that the latent space is sufficiently robust without strong KL regularization.
@@
D.2 Additional details of comparison across generative paradigms
@@
We also present extended results for our conditional generation tasks using a broader variety of evaluation metrics. For XSum and SQuAD2.0, we report ROUGE scores [16], a standard metric that evaluates text quality based on n-gram overlap with reference outputs. For ParaDetox, we additionally include the J-score, which is defined as the product of style accuracy, fluency, and content preservation. Extended results are summarized in Table˜9.
Appendix E Limitations
@@
Appendix F Societal Impact
@@
Appendix G Generation examples